the perimeter of a rectangle is 42 meters the lengh of the rectangle is 3 meters less than twice the width ,find the length and width of the rectangle
Found 2 solutions by checkley77, stanbon:
Answer by checkley77(12844)

L=2W-3
P=2L+2W
42=2(2W-3)+2W
42=4W-6+2W
6W=42+6
6W=48
W=48/6
W=8 ANS. FOR THE WIDTH.
L=2*8-3
L=16-3=13 ANS. FOR THE LENGTH.
PROOF:
42=2*13+2*8
42=26+16
42=42

Answer by stanbon(75887)

The perimeter of a rectangle is 42 meters. The length of the rectangle is 3 meters less than twice the width. Find the length and width of the rectangle.
----------------
P = 2(W + L)
42 = 2(W + (2W-3))
21 = 3W - 3
24 = 3W
W = 8 meters
===============
Cheers,
Stan H.
===============